# Surgery and positive Bakry–Émery Ricci curvature

### v1

@@ 1. Introduction
This result was independently also obtained by Ketterer [24], who additionally proved a converse in terms of the curvature-dimension condition $\mathrm{CD}(K,N)$ (see [24, Theorem 1.4]). We note that our proof of Theorem A, which is based on Perelman’s work [33], differs from Ketterer’s proof, which is based on a construction of Kosovskii [26], see also [40].
@@ Acknowledgements.
F.T. would like to thank the Centro di Ricerca Matematica Ennio De Giorgi and the Scuola Normale Superiore for their hospitality and support. She also acknowledges the support and hospitality of the Department of Mathematics of Fribourg during her visit. P.R. would like to thank Christian Ketterer and David Wraith for helpful conversations. Finally, both authors would like to thank David Wraith for his suggestions on an earlier version of this article.
@@ 2.2. Weighted Riemannian metrics on a cylinder
These functions are constructed in [35, Sections 3.2 and 3.3]. Notice that, compared to the corresponding statement in [35], we changed the notation and denoted $h$, $f$, $\lambda$ and $\cos(R/N)$ in [35] by $\beta$, $\gamma$, $\varepsilon$ and $\lambda$, respectively. We also omitted conclusions (3.7) and (3.10) being positive in [35], as we do not need them here.
@@ 3. Perelman’s Gluing Theorem
The construction of the metric $\tilde{h}$ via a cubic spline interpolation is carried out in [5, Lemma 3 and subsequent paragraph], see also [38, Section 2.1]. We are therefore left with constructing the function $\tilde{f}$. Here we take a similar approach and define $\tilde{f}$ as follows. Let us denote by $f_{i}(t)=f_{i}(t,\cdot)$ the induced weight function of $f_{i}$ on the hypersurface $\{t\}\times X$ at constant distance $t\in[-\varepsilon,\varepsilon]$ from $X$, with $i=1$ for $t\leq 0$ and $i=2$ for $t\geq 0$. We then get our desired weight function $\tilde{f}(t):=\tilde{f}(t,\cdot)$ via a cubic spline interpolation between $f_{1}(-\varepsilon)$ and $f_{2}(\varepsilon)$, that is
@@
The statement follows directly once we apply the smoothing result of [38, Lemma 3.1] to the $C^{1}$-regular weight function and the metric of the glued manifold $(M\cup_{\phi}M_{2},\tilde{h},e^{-\tilde{f}})$ constructed in Proposition 3.4. ∎
@@ 4. Connected Sums
We note that this definition slightly differs from the definition introduced in [10]. However, a core metric in the sense of Definition 4.1 can always be deformed into a core metric in the sense of [10] and vice versa, see [36, Lemma 2.4].
@@
An immediate consequence is that a closed manifold with a core metric is simply-connected. This follows from Theorem 4.2 in combination with the theorem of Bonnet–Myers, or alternatively from a result of Lawson [28, Theorem 1], which only requires that the mean curvature of $\varphi(S^{n-1})$ is non-negative.
@@
The proof goes along the same lines as the proof of [36, Lemma 2.18] by using Theorem A instead of Theorem 3.1. ∎
It is clear that a core metric defines a weighted core metric with respect to any $q\in(0,\infty]$ by choosing a constant weight function. Furthermore, just like for core metrics, closed manifolds with a weighted core metric are simply-connected. This follows from a result of Moore–Woolgar [31, Theorem 1.5], which only requires the weighted mean curvature on the boundary be non-negative, or, alternatively, from Theorem B in combination with Proposition A.2. A further connection between core metrics and weighted core metrics is given in Proposition A.6 below.
@@
It now remains to prove Proposition 4.4. The first step in the proof is to deform the weighted metric locally around the point $x$ with $\nabla f_{x}=0$ into a weighted metric with constant weight function and constant sectional curvature equal to $1$. This is a direct adaptation of the corresponding statement for positive Ricci curvature first established in [18] (where it is shown for negative Ricci curvature, but the arguments work similarly for positive Ricci curvature), see also [50, Theorem 1.10] and [39, Lemma 4.3]. We follow the line of arguments given in [39, Lemma 4.3 and Corollary 4.4] and begin by establishing the following more general deformation result first.
@@ 5. Higher Surgeries
To simplify notation, we will write $\varphi(S^{a})$ instead of $\varphi(S^{p}\times\{0\})$. Let $g_{r,\varepsilon}$ denote the metric $g_{\varphi(S^{a}),r}$ on $D_{\varepsilon}\nu(\varphi(S^{a}))$ considered above. For $r$ sufficiently small, this metric has positive Ricci curvature, see e.g. [4, 9.70]. Following Lemma 5.2, we view $g_{r,\varepsilon}$ as a metric on a tubular neighbourhood of $\varphi(S^{a})$. Therefore, Lemma 5.2 implies that the $1$-jets of the weighted metrics $(g,e^{-f})$ and $(g_{r,\varepsilon},e^{-f_{0}})$, where $f_{0}$ is the constant value of $f$ on $\varphi(S^{a})$, coincide on $\varphi(S^{a})$. Using weighted metrics of $\widetilde{\mathrm{Ric}}_{q}>0$, one can then deform $(g,e^{-f})$ into a new weighted metric $(\bar{g},e^{-\bar{f}})$ that coincides with $(g_{r,\varepsilon},e^{-f_{0}})$ on a neighbourhood of $\varphi(S^{a})$, and with $(g,e^{-f})$ outside a (bigger) neighbourhood of $\varphi(S^{a})$. We stress that these neighbourhoods can be chosen arbitrarily small, see [2, Theorem 1.2] and [50, Theorem 1.10].
@@ 6.1. Wall’s classification of handlebodies
We can determine the associated extended quadratic form of a handlebody from the attaching maps as follows. Let $\varphi_{1},\dots,\varphi_{\ell}\colon S^{2m}\times D^{2m+1}\hookrightarrow S^{4m+1}$ be embeddings with pairwise disjoint images. We extend each embedding $\varphi_{i}|_{S^{2m}\times\{0\}}$ to a map $D^{2m+1}\hookrightarrow D^{4m+2}$. By the relative Whitney embedding theorem [47, Theorem 5], we can assume that these maps are embeddings, and by the uniqueness of tubular neighbourhoods (see e.g. [20, Section 4.5]) we can extend them to embeddings $\bar{\varphi}_{i}\colon D^{2m+1}\times D^{2m+1}\hookrightarrow D^{4m+2}$ satisfying
@@ 6.2. Geometric setup
Finally, we define $h_{\varepsilon}$ as the function obtained from $\tilde{h}_{\varepsilon}$ by smoothing at the points of non-smoothness. For that, around each of these points, we choose neighbourhoods of size so small that they do not contain the points $t=\varepsilon,3\varepsilon,6\varepsilon$, and smooth out the function $\tilde{h}_{\varepsilon}$ using Corollary 3.6 (where we use a constant weight function), see also [35, Corollary 3.2]. To guarantee that the Ricci curvatures are still positive for the resulting function, we need to verify that at each of these points the left-hand side derivative is at least the right-hand side derivative. This is clear at the points $t=\nu\varepsilon,t_{\varepsilon},3\nu\varepsilon$. At $t=\frac{6}{\nu}\varepsilon$ this follows from the fact that the quotient $\frac{h_{2}^{\prime}(\frac{6}{\nu}\varepsilon)}{\sin(\frac{6}{\nu}\varepsilon)}$ converges to a value strictly bigger than $-1$ for $\nu$ sufficiently close to $1$ as seen above. ∎
@@ 6.4. Intersection matrices
Furthermore, by Lemmas 6.10 and 6.12 and Proposition 6.7, there exists a metric of positive Ricci curvature on $S^{4m+1}$ and embeddings $\bar{\varphi}_{i}\colon D^{2m+1}\hookrightarrow D^{4m+2}$, $1\leq i\leq 2\ell k$, with oriented intersection matrix $B_{\nu,\ell}$ and such that each restriction $\bar{\varphi}_{i}|_{S^{2m}}\colon S^{2m}\hookrightarrow S^{4m+1}$ is round and totally geodesic. Moreover, the embeddings $\bar{\varphi}_{i}|_{S^{2m}}$ have pairwise disjoint image. We extend each embedding $\bar{\varphi}_{i}|_{S^{2m}}$ to an embedding $\varphi_{i}\colon S^{2m}\times D^{2m+1}\hookrightarrow S^{4m+1}$ such that the invariants of the handlebody obtained from $\varphi_{1},\dots,\varphi_{\ell}$ are given by
@@ 6.6. Simply-connected 5-manifolds
Items (1) and (2) are shown in [29, Theorem 1] and [17, Corollary 4.4], respectively. The last statement follows from the fact that any closed, simply-connected manifold of dimension at least $5$ admits a metric of positive scalar curvature if and only if it is non-spin [19] or spin with vanishing $\alpha$-invariant [44], and the only closed, simply-connected manifolds in dimensions $2$ and $3$ are spheres. ∎
@@
Given the result of Proposition A.2, it is not clear, however, how one can construct a metric of positive scalar curvature from a weighted metric with positive weighted Ricci curvature. Note that in the special case where the weighted Ricci curvature is constant, the metric itself already has positive scalar curvature by [13, Proposition 1.1]. In general, however, one can construct examples, where the metric $g$ even has negative sectional curvature, see e.g. [46, Example 2.2]. On the other hand, for closed manifolds, there exists at least a point of positive Ricci curvature:
@@
This follows from the fact that $M\times S^{p}$ admits a submersion metric of $\mathrm{Ric}>0$ for all $p\geq\max\{2,q\}$ by [29, Section 2] or Proposition A.5 below (where we can replace $q$ by a finite value in case $q=\infty$ since $M$ is compact), together with [34, Theorem 2].
@@
For this theorem, one needs to extend the definition of $\widetilde{\mathrm{Ric}}_{q}$ to $q<0$ in the obvious way. We also note that in [53, Section 5.1] several examples are given where the holonomy of $\nabla^{df}$ is not compact, showing that Theorem A.4 does not provide a full answer to Question A.1.
As a variation of Question A.1, we can ask whether the existence of a weighted Riemannian metric of $\widetilde{\mathrm{Ric}}_{q}>0$ on a manifold $M$ implies the existence of a Riemannian metric of $\mathrm{Ric}>0$ on some higher-dimensional manifold obtained from $M$, such as a fibre bundle with base $M$. It was observed by Lott [29, Section 2] that this holds in the special case of a product $M\times S^{p}$ whenever $p\geq\max\{2,q\}$, i.e. if $M$ admits a weighted Riemannian metric of $\widetilde{\mathrm{Ric}}_{q}>0$, then $M\times S^{p}$ admits a Riemannian metric of $\mathrm{Ric}>0$ whenever $p\geq\max\{2,q\}$. We can generalise this as follows.
@@
In particular, the assumptions of Proposition A.6 are satisfied when $\pi$ is the trivial bundle $M\times N\to M$ and $N$ admits a core metric (then we can set $\hat{g}=\hat{g}^{\prime}$), or when $\pi$ is a linear sphere bundle (then we can set $\hat{g}=\hat{g}^{\prime}=ds_{p}^{2}$). Further, it can be applied to projective bundles with fibre $\mathbb{C}P^{n}$, $\mathbb{H}P^{n}$ or $\mathbb{O}P^{2}$, see [36, Section 5.2].
@@
We now consider the manifold $E\setminus\pi^{-1}(\varphi(D^{p}))^{\circ}$ equipped with the induced metric. By [9, Proposition 1.2.11] we can deform the metric $\bar{g}_{t}$ preserving $\mathrm{Ric}>0$ so that the second fundamental form on the boundary is strictly positive, and for any $r>0$, by [11, Theorem C], we can assume that the metric on the boundary is given by $ds_{p-1}^{2}+r^{2}\hat{g}^{\prime}$. Then, by [36, Theorem 4.1], it follows that we can glue back in $D^{p}\times N$ and obtain a core metric on $E$. ∎

### v2

@@ 1. Introduction
This result was independently also obtained by Ketterer [25], who additionally proved a converse in terms of the curvature-dimension condition $\mathrm{CD}(K,N)$ (see [25, Theorem 1.4]). We note that our proof of Theorem A, which is based on Perelman’s work [34], differs from Ketterer’s proof, which is based on a construction of Kosovskii [27], see also [41].
@@
As pointed out by Erik Hupp, the case where $q=2$ and all $M_{i}$ with $i\geq 1$ are given by $\mathbb{C}P^{2}$ in Theorem B also follows from the construction in [21].
@@ Acknowledgements.
F.T. would like to thank the Centro di Ricerca Matematica Ennio De Giorgi and the Scuola Normale Superiore for their hospitality and support. She also acknowledges the support and hospitality of the Department of Mathematics of Fribourg during her visit. P.R. would like to thank Christian Ketterer and David Wraith for helpful conversations. Finally, both authors would like to thank David Wraith for his suggestions on an earlier version of this article, Manuel Krannich for his explanations on smoothing theory, and Erik Hupp for making us aware of the article [21].
@@ 2.2. Weighted Riemannian metrics on a cylinder
These functions are constructed in [36, Sections 3.2 and 3.3]. Notice that, compared to the corresponding statement in [36], we changed the notation and denoted $h$, $f$, $\lambda$ and $\cos(R/N)$ in [36] by $\beta$, $\gamma$, $\varepsilon$ and $\lambda$, respectively. We also omitted conclusions (3.7) and (3.10) being positive in [36], as we do not need them here.
@@ 3. Perelman’s Gluing Theorem
The construction of the metric $\tilde{h}$ via a cubic spline interpolation is carried out in [5, Lemma 3 and subsequent paragraph], see also [39, Section 2.1]. We are therefore left with constructing the function $\tilde{f}$. Here we take a similar approach and define $\tilde{f}$ as follows. Let us denote by $f_{i}(t)=f_{i}(t,\cdot)$ the induced weight function of $f_{i}$ on the hypersurface $\{t\}\times X$ at constant distance $t\in[-\varepsilon,\varepsilon]$ from $X$, with $i=1$ for $t\leq 0$ and $i=2$ for $t\geq 0$. We then get our desired weight function $\tilde{f}(t):=\tilde{f}(t,\cdot)$ via a cubic spline interpolation between $f_{1}(-\varepsilon)$ and $f_{2}(\varepsilon)$, that is
@@
The statement follows directly once we apply the smoothing result of [39, Lemma 3.1] to the $C^{1}$-regular weight function and the metric of the glued manifold $(M\cup_{\phi}M_{2},\tilde{h},e^{-\tilde{f}})$ constructed in Proposition 3.4. ∎
@@ 4. Connected Sums
We note that this definition slightly differs from the definition introduced in [10]. However, a core metric in the sense of Definition 4.1 can always be deformed into a core metric in the sense of [10] and vice versa, see [37, Lemma 2.4].
@@
An immediate consequence is that a closed manifold with a core metric is simply-connected. This follows from Theorem 4.2 in combination with the theorem of Bonnet–Myers, or alternatively from a result of Lawson [29, Theorem 1], which only requires that the mean curvature of $\varphi(S^{n-1})$ is non-negative.
@@
The proof goes along the same lines as the proof of [37, Lemma 2.18] by using Theorem A instead of Theorem 3.1. ∎
It is clear that a core metric defines a weighted core metric with respect to any $q\in(0,\infty]$ by choosing a constant weight function. Furthermore, just like for core metrics, closed manifolds with a weighted core metric are simply-connected. This follows from a result of Moore–Woolgar [32, Theorem 1.5], which only requires the weighted mean curvature on the boundary be non-negative, or, alternatively, from Theorem B in combination with Proposition A.2. A further connection between core metrics and weighted core metrics is given in Proposition A.6 below.
@@
It now remains to prove Proposition 4.4. The first step in the proof is to deform the weighted metric locally around the point $x$ with $\nabla f_{x}=0$ into a weighted metric with constant weight function and constant sectional curvature equal to $1$. This is a direct adaptation of the corresponding statement for positive Ricci curvature first established in [18] (where it is shown for negative Ricci curvature, but the arguments work similarly for positive Ricci curvature), see also [51, Theorem 1.10] and [40, Lemma 4.3]. We follow the line of arguments given in [40, Lemma 4.3 and Corollary 4.4] and begin by establishing the following more general deformation result first.
@@ 5. Higher Surgeries
To simplify notation, we will write $\varphi(S^{a})$ instead of $\varphi(S^{p}\times\{0\})$. Let $g_{r,\varepsilon}$ denote the metric $g_{\varphi(S^{a}),r}$ on $D_{\varepsilon}\nu(\varphi(S^{a}))$ considered above. For $r$ sufficiently small, this metric has positive Ricci curvature, see e.g. [4, 9.70]. Following Lemma 5.2, we view $g_{r,\varepsilon}$ as a metric on a tubular neighbourhood of $\varphi(S^{a})$. Therefore, Lemma 5.2 implies that the $1$-jets of the weighted metrics $(g,e^{-f})$ and $(g_{r,\varepsilon},e^{-f_{0}})$, where $f_{0}$ is the constant value of $f$ on $\varphi(S^{a})$, coincide on $\varphi(S^{a})$. Using weighted metrics of $\widetilde{\mathrm{Ric}}_{q}>0$, one can then deform $(g,e^{-f})$ into a new weighted metric $(\bar{g},e^{-\bar{f}})$ that coincides with $(g_{r,\varepsilon},e^{-f_{0}})$ on a neighbourhood of $\varphi(S^{a})$, and with $(g,e^{-f})$ outside a (bigger) neighbourhood of $\varphi(S^{a})$. We stress that these neighbourhoods can be chosen arbitrarily small, see [2, Theorem 1.2] and [51, Theorem 1.10].
@@ 6. Highly-connected manifolds
Remark 6.2.
Similarly as in Corollary E, it follows from Theorem 6.1 that for any such manifold $M$ there exists a homotopy sphere $\Sigma^{4m+1}$ such that for all $q\in\mathbb{N}$ sufficiently large, the manifold $(M\#\Sigma)\times S^{q}$ admits a Riemannian metric of $\mathrm{Ric}>0$. It remains an open question whether the connected sum with $\Sigma$ is necessary, i.e. whether the manifold $M\times S^{q}$ admits a Riemannian metric of $\mathrm{Ric}>0$. This would follow directly if one could show that $(M\#\Sigma)\times S^{q}$ is in fact diffeomorphic to $M\times S^{q}$.
If we identify $M\times S^{q}$ and $(M\#\Sigma)\times S^{q}$ as topological manifolds in the obvious way, easy arguments in smoothing theory show that the corresponding smooth structures are not concordant, and hence not isotopic, when $\Sigma$ is not the standard sphere. However, this does not rule out the existence of a diffeomorphism between the two smooth structures. To the best of our knowledge, this problem is open.
@@ 6.1. Wall’s classification of handlebodies
We can determine the associated extended quadratic form of a handlebody from the attaching maps as follows. Let $\varphi_{1},\dots,\varphi_{\ell}\colon S^{2m}\times D^{2m+1}\hookrightarrow S^{4m+1}$ be embeddings with pairwise disjoint images. We extend each embedding $\varphi_{i}|_{S^{2m}\times\{0\}}$ to a map $D^{2m+1}\hookrightarrow D^{4m+2}$. By the relative Whitney embedding theorem [48, Theorem 5], we can assume that these maps are embeddings, and by the uniqueness of tubular neighbourhoods (see e.g. [20, Section 4.5]) we can extend them to embeddings $\bar{\varphi}_{i}\colon D^{2m+1}\times D^{2m+1}\hookrightarrow D^{4m+2}$ satisfying
@@ 6.2. Geometric setup
Finally, we define $h_{\varepsilon}$ as the function obtained from $\tilde{h}_{\varepsilon}$ by smoothing at the points of non-smoothness. For that, around each of these points, we choose neighbourhoods of size so small that they do not contain the points $t=\varepsilon,3\varepsilon,6\varepsilon$, and smooth out the function $\tilde{h}_{\varepsilon}$ using Corollary 3.6 (where we use a constant weight function), see also [36, Corollary 3.2]. To guarantee that the Ricci curvatures are still positive for the resulting function, we need to verify that at each of these points the left-hand side derivative is at least the right-hand side derivative. This is clear at the points $t=\nu\varepsilon,t_{\varepsilon},3\nu\varepsilon$. At $t=\frac{6}{\nu}\varepsilon$ this follows from the fact that the quotient $\frac{h_{2}^{\prime}(\frac{6}{\nu}\varepsilon)}{\sin(\frac{6}{\nu}\varepsilon)}$ converges to a value strictly bigger than $-1$ for $\nu$ sufficiently close to $1$ as seen above. ∎
@@ 6.4. Intersection matrices
Furthermore, by Lemmas 6.11 and 6.13 and Proposition 6.8, there exists a metric of positive Ricci curvature on $S^{4m+1}$ and embeddings $\bar{\varphi}_{i}\colon D^{2m+1}\hookrightarrow D^{4m+2}$, $1\leq i\leq 2\ell k$, with oriented intersection matrix $B_{\nu,\ell}$ and such that each restriction $\bar{\varphi}_{i}|_{S^{2m}}\colon S^{2m}\hookrightarrow S^{4m+1}$ is round and totally geodesic. Moreover, the embeddings $\bar{\varphi}_{i}|_{S^{2m}}$ have pairwise disjoint image. We extend each embedding $\bar{\varphi}_{i}|_{S^{2m}}$ to an embedding $\varphi_{i}\colon S^{2m}\times D^{2m+1}\hookrightarrow S^{4m+1}$ such that the invariants of the handlebody obtained from $\varphi_{1},\dots,\varphi_{\ell}$ are given by
@@ 6.6. Simply-connected 5-manifolds
Items (1) and (2) are shown in [30, Theorem 1] and [17, Corollary 4.4], respectively. The last statement follows from the fact that any closed, simply-connected manifold of dimension at least $5$ admits a metric of positive scalar curvature if and only if it is non-spin [19] or spin with vanishing $\alpha$-invariant [45], and the only closed, simply-connected manifolds in dimensions $2$ and $3$ are spheres. ∎
@@
Given the result of Proposition A.2, it is not clear, however, how one can construct a metric of positive scalar curvature from a weighted metric with positive weighted Ricci curvature. Note that in the special case where the weighted Ricci curvature is constant, the metric itself already has positive scalar curvature by [13, Proposition 1.1]. In general, however, one can construct examples, where the metric $g$ even has negative sectional curvature, see e.g. [47, Example 2.2]. On the other hand, for closed manifolds, there exists at least a point of positive Ricci curvature:
@@
This follows from the fact that $M\times S^{p}$ admits a submersion metric of $\mathrm{Ric}>0$ for all $p\geq\max\{2,q\}$ by [30, Section 2] or Proposition A.5 below (where we can replace $q$ by a finite value in case $q=\infty$ since $M$ is compact), together with [35, Theorem 2].
@@
For this theorem, one needs to extend the definition of $\widetilde{\mathrm{Ric}}_{q}$ to $q<0$ in the obvious way. We also note that in [54, Section 5.1] several examples are given where the holonomy of $\nabla^{df}$ is not compact, showing that Theorem A.4 does not provide a full answer to Question A.1.
As a variation of Question A.1, we can ask whether the existence of a weighted Riemannian metric of $\widetilde{\mathrm{Ric}}_{q}>0$ on a manifold $M$ implies the existence of a Riemannian metric of $\mathrm{Ric}>0$ on some higher-dimensional manifold obtained from $M$, such as a fibre bundle with base $M$. It was observed by Lott [30, Section 2] that this holds in the special case of a product $M\times S^{p}$ whenever $p\geq\max\{2,q\}$, i.e. if $M$ admits a weighted Riemannian metric of $\widetilde{\mathrm{Ric}}_{q}>0$, then $M\times S^{p}$ admits a Riemannian metric of $\mathrm{Ric}>0$ whenever $p\geq\max\{2,q\}$. We can generalise this as follows.
@@
In particular, the assumptions of Proposition A.6 are satisfied when $\pi$ is the trivial bundle $M\times N\to M$ and $N$ admits a core metric (then we can set $\hat{g}=\hat{g}^{\prime}$), or when $\pi$ is a linear sphere bundle (then we can set $\hat{g}=\hat{g}^{\prime}=ds_{p}^{2}$). Further, it can be applied to projective bundles with fibre $\mathbb{C}P^{n}$, $\mathbb{H}P^{n}$ or $\mathbb{O}P^{2}$, see [37, Section 5.2].
@@
We now consider the manifold $E\setminus\pi^{-1}(\varphi(D^{p}))^{\circ}$ equipped with the induced metric. By [9, Proposition 1.2.11] we can deform the metric $\bar{g}_{t}$ preserving $\mathrm{Ric}>0$ so that the second fundamental form on the boundary is strictly positive, and for any $r>0$, by [11, Theorem C], we can assume that the metric on the boundary is given by $ds_{p-1}^{2}+r^{2}\hat{g}^{\prime}$. Then, by [37, Theorem 4.1], it follows that we can glue back in $D^{p}\times N$ and obtain a core metric on $E$. ∎